Hi I would like to receive help solving for "Y" in terms of "X" for the problem below.
3X-5Y=25

Found 2 solutions by ewatrrr, rothauserc:
Answer by ewatrrr(24785)

3X-5Y=25
3X - 25 = 5Y
(3/5)X - 5 = Y
0R
Y = (3/5)X - 5

Answer by rothauserc(4718)

3X-5Y=25
subtract 3x from both sides of =
-5y = 25 - 3x
divide both sides of = by -5
y = -5 + 3x/5